A car of mass m is on an icy driveway inclined at an angle θ as in Figures below. Find the acceleration of the car, assuming that the driveway is frictionles
1
Expert's answer
2020-02-10T09:33:00-0500

Mass of a car: "m"

Angle: "\theta"

"\overline a" - ?

Solution:

"\overline{F}=m\overline g" , "\overline n"


"\overline F+\overline n=m\overline a"

"x: F_x+n_x=mg\sin \theta +0=ma_x"

"y:F_y+n_y=0"


"mg\sin\theta=ma_x, \quad 0=a_y"

"a_x=g\sin\theta=a"


Answer: "a=g\sin\theta" .